OpenStudy (anonymous):

3sin(3pix)+3sqrt^(3)cos(3pix)

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

Sorry! I forgot to clarify the question. "Write the expression in terms of sine only" 3sin(3pix)+3sqrt^(3)cos(3pix)

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

No problem, I can help now I know what you're asking. Could you clarify the expression a bit please? A screenshot/photo of the question would help. I don't understand what you mean by sqrt^ as ^ tends to denote an index but it doesn't make sense next to a sqrt.

OpenStudy (anonymous):

thank you so much!! @tom982

OpenStudy (anonymous):

Ah I see, for future reference that's usually denoted 3sqrt(3) :) I can't expand those in my head so I'm going to have to write it all down, bear with me.

OpenStudy (anonymous):

Ok thank you!! Sorry jeje

OpenStudy (anonymous):

Do you have possible answers? I got half way through working it out properly then realised there's a way that's kind of cheating, but technically speaking it isn't wrong.

OpenStudy (anonymous):

No :/ they don't give us any possible answers

OpenStudy (anonymous):

Well using the trig identity: \[\cos^2(x) + \sin^2(x) = 1\]We know that \[\cos^2(x) = 1 - sin^2(x)\] and therefore \[\cos(x) = \sqrt{1 - \sin^2(x)}\] So we can rewrite our equation as \[3\sin(4\pi x) + 3\sqrt{3}(1-\sin^2(4\pi x))\] The above equation is now in terms of sine only :)

OpenStudy (anonymous):

I got it! thank you so so much!!!

OpenStudy (anonymous):

No problem :)